# Funding AI for Good: A Call for Meaningful Engagement

### v1

@@ Abstract.
Artificial Intelligence for Social Good (AI4SG) is a growing area exploring AI’s potential to address social issues like public health. Yet prior work has shown limited evidence of its tangible benefits for intended communities, and projects frequently face inadequate community engagement and sustainability challenges. Funding agendas play a crucial role in framing AI4SG initiatives and shaping their approaches. Through a qualitative analysis of 35 funding documents—representing about \$410 million USD in total investments, we reveal dissonances between AI4SG’s stated intentions for positive social impact and the techno-centric approaches that some funding agendas promoted. Drawing on our findings, we offer recommendations for funders to scaffold approaches that balance both contextual understanding and technical capacities in future funding call designs. We call for greater engagement between AI4SG funders and the HCI community to support community engagement work in the funding program design process.
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXprice: 15.00isbn: XXXXX
@@ 1. Introduction
Figure 1. AI4SG aims to bring about positive, long-lasting social impact. Prior literature emphasizes that impactful innovations are the result of a balance of technology capacities and contextual understanding (right panel). A bias towards technology capacities leads to innovations without a real-world positive impact (left panel). We use this spectrum as a lens to investigate whether current AI4SG funding agendas orient towards more techno-centric or balanced approaches.
@@
Artificial Intelligence (AI) for social good (AI4SG) initiatives seek to bring about positive and long-lasting impact in social domains such as public health, by leveraging data-driven technologies like machine learning (Shi et al., 2020; Li et al., 2021; Perrault et al., 2020; Rolnick et al., 2022; Tomašev et al., 2020). However, a growing body of evidence suggests that AI4SG initiatives do not necessarily yield the intended positive outcomes in real-world contexts (Vinuesa et al., 2020; Eubanks, 2018; SALLY HO and GARANCE BURKE, 2023; Tate Ryan-Mosley, 2023). Research in Computer-Supported Cooperative Work (CSCW), Information and Communication Technologies and Development (ICTD), and Human-Computer Interaction for Development (HCI4D) provides insights into the socio-technical mechanisms underpinning challenges in delivering the tangible positive impact of AI. While existing work has largely focused on the roles and perspectives of project teams (often consisting of AI technologists and mission-driven organizations), emerging works point to the significant influence of funders in shaping the direction, design, and priorities of technology initiatives aimed at social impact (Lin et al., 2024; Saha et al., 2022a; Erete et al., 2023).
Funders are organizations that provide monetary support and facilitate interdisciplinary programs, such as private foundations, nonprofit organizations, and technology companies. Funders function not only as financial sponsors but also as agenda-setters who significantly shape the ecosystems of AI development (Widder et al., 2024; Gardner et al., 2022). In AI4SG specifically,  Lin et al. (2024) draw on community organizations’ perspectives and reveal funders’ prevalent influence in how AI4SG partnerships are formed, which problem spaces are prioritized, and what kinds of solutions are pursued during the early stages of projects. Prior work also points out that funding often incentivizes mission-driven organizations’ tendency to adopt data-driven innovations, even at the cost of their original social missions (Bopp et al., 2017; Lin et al., 2024). For AI developers and practitioners, short funding timelines often constrain their considerations of ethics and unintended consequences (Do et al., 2023; Moitra et al., 2022; Okolo and Lin, 2024). While prior work has established funders’ prevalent influence through the perspectives of mission-driven organizations and AI developers in AI4SG, we still lack a direct understanding of how funders frame AI’s role in social good problems and how they envision AI’s social impact might be achieved.
To fill this gap, we take up Laura Nader’s (1972) call to “study up” AI4SG by analyzing its funding agendas (Nader, 1972). By focusing on funding, our work seeks to draw attention away from assigning blame to AI4SG project teams, and towards recognizing the broader power structures they are embedded in. Since AI4SG builds upon prior “tech for good” efforts (Aula and Bowles, 2023), we specifically draw on lessons learned from CSCW, ICTD, and HCI4D. This prior literature has extensively examined the approaches that influence the outcomes of socially driven technology projects. It emphasizes that in order for innovations to be impactful in real-world contexts, they must be grounded in both a thorough understanding of the problem and access to appropriate technological capabilities (Figure 1). A spectrum from techno-centric to balanced approaches emerges from our synthesis of this rich body of work, which points to a strong association between where projects fall on this spectrum and project outcomes. A techno-centric approach assumes novel technology alone can solve complex social issues, often leading to suboptimal results and strained relationships between researchers and communities (e.g., (Ames, 2019; Karusala and Kumar, 2017; Walsham, 2017; Dada, 2006; Brown and Mickelson, 2019)). In contrast, a more balanced approach is grounded in a solid understanding of the community’s challenges and contexts through community-collaborative approaches (Cooper and Zafiroglu, 2024), and tends to result in more community buy-in and sustainable efforts (e.g., (Toyama, 2015; Gandhi et al., 2007; Burrell and Toyama, 2009)). Through the lens of this spectrum, our study investigates whether current AI4SG funding agendas orient towards a more techno-centric or a more balanced approach. Specifically, we ask the following research questions:
@@
  • RQ1: How do funders frame social good, and what is AI’s envisioned role in advancing social good, as articulated in funding documents?
  • RQ2: What are funders’ expected outcomes of AI4SG projects?
  • RQ3: How do funders envision these goals might be achieved?
@@
To address these questions, we conducted a qualitative analysis of AI4SG funding calls for proposals and grant announcements (hereafter referred to as funding documents). Our dataset encompasses 35 funding documents—19 funding calls and 16 grant announcements—representing about \$410 million USD in total investments (see Appendix A for a complete list). These programs span a range of efforts, broadly framed under the banners of AI or data science for “social good,” “development,” “climate change,” and other social domains. Our dataset includes funding programs from 23 distinct funders, including private technology companies, philanthropic foundations, nonprofit organizations, and government agencies based in English-speaking countries such as the United States.
@@
Funding documents in our corpus overwhelmingly mobilized AI against large-scale social problems that cannot be fully addressed with any single intervention (notably public health and development). Yet several funding documents framed AI’s potential as deterministically beneficial and positioned AI interventions as solutions to these social problems while skirting contextual and stakeholder specificity. Some others scoped interventions to more well-defined sub-problems and explicitly acknowledged both AI’s benefits and risks. Expected outcomes of funding programs in our dataset frequently centered on technical deliverables, while a subset adopted broader success metrics attentive to benefits to the communities involved. Furthermore, although interdisciplinary collaboration was widely encouraged, funding documents varied in their emphasis and scaffolding of community engagement and who were considered “experts.” With a few exceptions, expectations for community engagement stopped at a consultation mode towards later stages of projects (e.g., implementation stage), where community stakeholders were expected to provide inputs but not able to shape the broader project scopes.
@@
To the best of our knowledge, this study presents the first qualitative analysis of how AI4SG funding documents frame social good, envision AI’s role in advancing social good, and anticipate desired outcomes to be achieved. While prior work has pointed out AI4SG’s tendency towards more techno-centric approaches, we provide detailed insights into how both techno-centric and balanced approaches are instantiated in different components of funding agendas. Contributing to existing literature on “tech for good” initiatives, we first highlight the disconnects between the positive social impact that AI4SG aims for and the techno-centric approaches that some funding documents promoted (Section 5.1). We then build on the more balanced approaches and existing gaps we identified in our dataset and propose additional opportunities for designing future funding calls, including supporting relationship-building, maintenance post-project deployment, and community engagement training for project teams (Section 5.2). Finally, we outline an agenda for meaningful community engagement for funders, emphasizing the role that the HCI community can and should play in helping shape AI4SG funding agendas and design processes (Section 5.3).
@@ 2. Related Work
Prior HCI research on AI4SG has highlighted both the movement’s intention for positive social impact and its challenges (Section 2.1). Scholars have called on AI4SG practitioners to draw on lessons from previous “tech for good” projects, which emphasize that impactful innovations require a balance of appropriate technical capacity and deep contextual understanding (Section 2.2). While prior work points out AI4SG’s bias towards emphasizing technical capacity, our work contributes an analysis of how this bias is instantiated across different components of funding documents, while offering concrete entry points for deeper contextual understanding. Despite the central role that funders play in shaping the AI4SG ecosystem, there is limited scholarship examining how funders frame AI4SG and what approaches are envisioned by funders (Section 2.3). In this study, we fill this gap by examining AI4SG funding documents through the lens of existing literature on “tech for good” projects.
@@ 2.1. The Promises and Challenges of AI for Social Good
Artificial Intelligence for Social Good (AI4SG) has emerged as a growing field of research and practice focused on applying AI to complex social issues (Shi et al., 2020; Perrault et al., 2020; Tomašev et al., 2020; Rolnick et al., 2022). Its stated goal is to deliver transformative, positive, and lasting impact across many social aspects (Perera, 2024; Cowls, 2021; Tudor et al., 2024). Some notable examples include projects in agriculture that predict crop yields using remote sensing data (You et al., 2017) and provide price forecasting for farmers in India (Ma et al., 2019). There is a growing interest in both use-inspired AI4SG research (Shi et al., 2020; IJCAI, 2024) and real-world implementations (e.g., (Data Science for Social Good Fellowship, 2024; ITU, 2024)).
@@
While enthusiasm for AI4SG remains high, growing evidence shows that many projects fall short of delivering lasting positive social impact and often end at the publication stage without deployment (Lin et al., 2024; Aula and Bowles, 2023; Ben Green, 2019). In high-stakes contexts such as public service screening or criminal justice, AI may be inappropriate altogether (Eubanks, 2018; SALLY HO and GARANCE BURKE, 2023; Pruss, 2023). Additionally, AI4SG is often shaped by promotional hype (Joyce and Cruz, 2024), magical thinking (Lupetti and Murray-Rust, 2024), and “enchanted determinism” (Campolo and Crawford, 2020), where the fascination with the mysterious capabilities of new technologies captures public attention toward techno-solutionism—where complex, structural challenges are reduced to problems amenable to AI interventions (Morozov, 2013). This could be exacerbated by AI’s transparency issues (especially in the Global South context that AI4SG projects are intended for (Okolo and Lin, 2024)). For instance, studies in India show a strong belief in “AI authority” over public institutions and lowered expectations for accountability (Kapania et al., 2022; Ramesh et al., 2022).
@@
HCI scholars have started to examine the mechanisms behind AI4SG initiatives. We highlight two key insights below. First, the aspiration to employ AI for social good is not a recent phenomenon; AI4SG should draw lessons learned from prior “tech for good” movements (Aula and Bowles, 2023; Lin et al., 2024). ICTD and HCI4D have long explored the development and use of emerging technologies in marginalized communities, especially in the Global South (Walsham, 2017; Dell and Kumar, 2016; Avgerou, 2010). Second, while past work often focuses on investigating perspectives of project teams—typically AI developers and nonprofit organizations, emerging research highlights funders’ role in shaping interdisciplinary partnerships (Lin et al., 2024; Saha et al., 2022a; Erete et al., 2023). Through interviews with staff of community organizations participating in AI4SG,  Lin et al. (2024) found that funding agenda influences program goals, proposed solutions, and definitions of success. Focusing on nonprofit organizations, other studies show how funding agendas shape their incentives to adopt emerging tech, and collaboration choices (Bopp et al., 2017; Erete et al., 2016; Goldenfein and Mann, 2023). AI developers and researchers’ approaches to AI ethics and considerations of unintended consequences are also shown to be impacted by funding timeline (Do et al., 2023; Widder et al., 2023; Moitra et al., 2022). Emphasizing the need to examine how power structures impact technology development in social contexts, scholars have taken up Laura Nader’s (1972) call to “study up” by focusing on those who shape social systems and exert normative influence (Nader, 1972). Emerging scholarship in algorithmic fairness (Barabas et al., 2020), public sector AI (Kawakami et al., 2024a), and machine learning data (Miceli et al., 2022) demonstrates that examining those in power can reveal critical leverage points for more accountable AI development.
While prior work has highlighted funders’ significant influence—primarily through the perspectives of mission-driven organizations and AI developers—we still lack a direct understanding of how funders frame AI’s role in addressing social good challenges. In this study, we “study up” AI4SG funding documents, through which funders communicate their intentions and envisioned approaches to AI4SG.

### v2

@@ Abstract.
Artificial Intelligence for Social Good (AI4SG) is a growing area that explores AI’s potential to address social issues, such as public health. Yet prior work has shown limited evidence of its tangible benefits for intended communities, and projects frequently face inadequate community engagement and sustainability challenges. While existing HCI literature on AI4SG initiatives primarily focuses on the mechanisms of funded projects and their outcomes, much less attention has been given to the funding agenda and rhetoric that influences downstream approaches. Through a thematic analysis of 35 funding documents—representing about \$410 million USD in total investments, we reveal dissonances between AI4SG’s stated intentions for positive social impact and the techno-centric approaches that some funding agendas promoted, while also identifying funding documents that scaffolded community-collaborative approaches for applicants. Drawing on our findings, we offer recommendations for funders to embed approaches that balance both contextual understanding and technical capacities in future funding call designs. We further discuss how the HCI community can positively shape AI4SG funding design processes.
copyright: acmcopyrightjournalyear: 2026doi: XXXXXXX.XXXXXXXprice: 15.00isbn: XXXXX
@@ 1. Introduction
Figure 1. AI4SG aims to bring about positive, long-lasting social impact. Prior literature emphasizes that impactful innovations are the result of a balance of technology capacities and contextual understanding (right panel). A bias towards technology capacities leads to innovations without a real-world positive impact (left panel). We use this spectrum as a lens to investigate how current AI4SG funding documents orient towards more techno-centric or balanced approaches.
@@
Artificial Intelligence (AI) for social good (AI4SG) initiatives aim to create long-lasting positive impact in social domains such as public health, by leveraging AI technologies like machine learning (Shi et al., 2020; Li et al., 2021; Perrault et al., 2020; Rolnick et al., 2022; Tomašev et al., 2020). Yet prior work finds limited evidence of its tangible benefits for intended communities, and projects frequently face deployment and sustainability challenges in real-world contexts (Vinuesa et al., 2020; Eubanks, 2018; SALLY HO and GARANCE BURKE, 2023; Tate Ryan-Mosley, 2023). HCI scholarship provides insights into the socio-technical mechanisms underpinning these challenges, and recent work points to the significant influence of funding agendas—the priorities, goals, and approaches set by funders—in shaping the technology initiatives aimed at social impact (Lin et al., 2024; Saha et al., 2022a; Erete et al., 2023). In AI4SG specifically,  Lin et al. (2024) show from community organizations’ perspectives that funding agendas heavily influence how partnerships are formed, which problem spaces are prioritized, and what kinds of solutions are pursued during the early stages of projects. Prior work also notes that funding often incentivizes mission-driven organizations to adopt data-driven innovations, sometimes at the cost of their original social missions (Bopp et al., 2017; Lin et al., 2024). For AI developers and practitioners, short funding timelines can constrain considerations of ethics and unintended consequences (Do et al., 2023; Moitra et al., 2022; Okolo and Lin, 2024). However, existing HCI literature on AI4SG initiatives primarily focuses on funded projects’ approaches and outcomes through the perspectives of project teams (typically consisting of AI technologists and mission-driven organizations); much less attention is given to the design of funding agendas themselves, before projects even get selected.
In this work, we begin addressing this gap by investigating AI4SG funding agendas, following up on Laura Nader’s call to “study up” (Nader, 1972) systems and actors that exert normative influence. In doing so, we aim to shift attention from assigning blame to AI4SG project teams, and towards recognizing the broader power structures shaping their approaches. We turn to public funding documents—textual instruments that communicate funding agendas, including calls for proposals and grant announcements. Prior literature has established that funding documents are not neutral instruments: they construct priorities (Palmero et al., 2025), shape applicant behaviors (Peng et al., 2024), and structure downstream research activities (Widder et al., 2024; Smith et al., 2023). In AI4SG, by defining who is eligible, what constitutes a “problem,” how these problems should be addressed, and what “success” means, funding documents construct discourses about AI and social good, signaling what approaches are fundable or imaginable. Although there is complexity in how written funding documents influence what actually gets funded, examining them reveals how priorities, values, and power relations are embedded even before any project begins. Funding documents can reinforce dominant AI imaginaries—“collective visions, beliefs, symbols, and expectations that individuals and communities hold about AI and its potential outcomes” (Zhong et al., 2025)—as seen in public texts portraying AI as “inevitable and massively disrupting” (Sophie Bennani-Taylor, 2024; Bareis and Katzenbach, 2022). Yet they can also guide applicants toward community engagement strategies previously unfamiliar to them.
In analyzing funding documents, our goal is not to evaluate how they eventually lead to funded project outcomes, but to provide thematic insights into the rhetoric and conceptual framing of AI4SG in funding documents that influence downstream approaches. Specifically, we ask the following research questions:
@@
  • RQ1: How do funding documents frame social good, and what is AI’s envisioned role in advancing social good, as articulated in funding documents?
  • RQ2: How do funding documents describe the expected outcomes of AI4SG projects?
  • RQ3: How do funding documents characterize how these goals might be achieved?
@@
We collected and conducted a reflexive thematic analysis of 35 funding documents—19 funding calls and 16 grant announcements—representing about \$410 million USD in total investments (see Appendix A for a complete list). These documents represent a range of efforts, broadly framed under the banners of AI or data-driven technologies for “social good,” “development,” and other social domains. Our dataset includes funding programs from 23 distinct funders, including private technology companies, philanthropic foundations, nonprofit organizations, and government agencies based in Western societies such as the United States.
Since AI4SG builds on earlier “tech for good” efforts (Aula and Bowles, 2023), we ground our analysis on lessons learned from Computer-Supported Cooperative Work (CSCW), Information and Communication Technologies and Development (ICTD), and Human-Computer Interaction for Development (HCI4D). This prior literature has extensively examined different approaches in socially driven technology projects, emphasizing that impactful innovations require both a thorough understanding of the problem and appropriate technological capabilities (Figure 1). From our synthesis of this body of work, a spectrum from techno-centric to balanced approaches emerges, indicating a strong association between where projects fall on this spectrum and project outcomes. A techno-centric approach assumes novel technology alone can solve complex social issues, often producing suboptimal results and strained relationships between researchers and communities (e.g., (Ames, 2019; Karusala and Kumar, 2017; Walsham, 2017; Dada, 2006; Brown and Mickelson, 2019)). In contrast, a more balanced approach, grounded in a solid understanding of the community’s challenges and contexts through community-collaborative approaches (Cooper and Zafiroglu, 2024), tend to result in more community buy-in and sustainability (e.g., (Toyama, 2015; Gandhi et al., 2007; Burrell and Toyama, 2009)). Using this spectrum, our study analyzes how current AI4SG funding rhetoric is oriented towards a more techno-centric or a more balanced approach.
@@
Funding documents in our corpus overwhelmingly mobilized AI against large-scale social problems that cannot be fully addressed with any single intervention (notably public health and development). Yet several funding documents framed AI’s potential as deterministically beneficial and positioned AI interventions as solutions to these social problems while skirting contextual and stakeholder specificity. Some other documents scoped interventions to more well-defined sub-problems and explicitly acknowledged both AI’s benefits and risks. Expected outcomes of funding documents in our dataset frequently centered on technical deliverables, while a subset adopted broader success metrics attentive to benefits to the communities involved. Furthermore, although interdisciplinary collaboration was widely encouraged, funding documents varied in their emphasis and scaffolding of community engagement and who were considered “experts.” With a few exceptions, expectations for community engagement stopped at a consultation mode towards later stages of projects (e.g., implementation stage), where community stakeholders were expected to provide inputs but not able to shape the broader project scopes.
To the best of our knowledge, this study presents the first qualitative analysis of how AI4SG funding documents frame social good, envision AI’s role in advancing it, and anticipate desired outcomes to be achieved. While prior work has largely focused on funded projects’ tendency towards more techno-centric approaches, we uncover how both techno-centric and balanced approaches are instantiated in different components of funding documents—before project selection even occurs. In doing so, this work makes three contributions to existing HCI literature on “tech for good” initiatives and document analysis of AI imaginaries:
@@
First, we provide a novel focus on AI4SG funding documents as a discourse instrument rather than neutral texts. Through a reflexive thematic analysis (Braun and Clarke, 2006, 2019), we highlight the disconnects between the positive social impact AI4SG aims for and the techno-centric approaches promoted in some funding documents (Section 5.1). Second, building on the more balanced approaches and gaps identified in our dataset, we propose opportunities for designing future funding calls for funders (including supporting relationship-building, maintenance post-project deployment, and community engagement training for project teams), and outline future research directions for HCI researchers (Section 5.2). Finally, we call for meaningful community engagement in the funding design process for funders, and specify the roles that HCI practitioners and researchers can play in shaping AI4SG funding agendas and design processes (Section 5.3).
@@ 2. Related Work
Prior HCI research on AI4SG has highlighted both the objective of existing funded projects to produce positive social impact and the challenges they face in doing so (Section 2.1). Scholars have called on AI4SG to draw on lessons from previous “tech for good” projects, which emphasize that impactful innovations require a balance of appropriate technical capacity and deep contextual understanding (Section 2.2). Given the central role of funding agendas in shaping the AI4SG ecosystem, we draw on emerging scholarship examining grant-making and private tech companies’ investment in AI4SG (Section 2.3). Furthermore, prior research in HCI and innovation studies has enacted document analysis to examine how AI is imagined in public texts and established the role of funding documents in shaping research priorities and collaboration structures (Section 2.4). However, HCI has yet to systematically analyze funding documents that define and construct AI4SG initiatives. This study begins to address this gap by analyzing AI4SG funding documents through the lens of existing literature on “tech for good.” While prior work points out funded projects’ bias towards emphasizing technical capacity, our work contributes an analysis of how this bias is instantiated across some funding documents, while also identifying more balanced approaches that emphasize contextual understanding.
@@ 2.1. The Promises and Challenges of AI for Social Good
Artificial Intelligence for Social Good (AI4SG) has emerged as a growing field of use-inspired research and real-world implementations focused on applying AI to complex social issues (Shi et al., 2020; Perrault et al., 2020; Tomašev et al., 2020; Rolnick et al., 2022). Its stated goal is to deliver transformative, positive, and lasting impact across many social aspects (Perera, 2024; Cowls, 2021; Tudor et al., 2024). Some notable examples include projects in agriculture that predict crop yields using remote sensing data (You et al., 2017) and provide price forecasting for farmers in India (Ma et al., 2019). While enthusiasm for AI4SG remains high, growing evidence shows that many projects fall short of delivering lasting positive social impact and often end at the publication stage without deployment (Lin et al., 2024; Aula and Bowles, 2023; Ben Green, 2019). In high-stakes contexts such as public service screening or criminal justice, AI may be inappropriate altogether (Eubanks, 2018; SALLY HO and GARANCE BURKE, 2023; Pruss, 2023). Additionally, AI4SG is often shaped by promotional hype (Joyce and Cruz, 2024), magical thinking (Lupetti and Murray-Rust, 2024), and “enchanted determinism” (Campolo and Crawford, 2020), where the fascination with the mysterious capabilities of new technologies captures public attention toward techno-solutionism—where complex, structural challenges are reduced to problems amenable to AI interventions (Morozov, 2013). This could be exacerbated by AI’s transparency issues (especially in the Global South context that AI4SG projects are intended for (Okolo and Lin, 2024)). For instance, studies in India show a strong belief in “AI authority” over public institutions and lowered expectations for accountability (Kapania et al., 2022; Ramesh et al., 2022).
@@
HCI scholars have started to examine the mechanisms behind AI4SG projects. We highlight two key insights below. First, the aspiration to employ AI for social good is not a recent phenomenon; AI4SG should draw lessons learned from prior “tech for good” movements (Aula and Bowles, 2023; Lin et al., 2024). ICTD and HCI4D have long explored the development and use of emerging technologies in marginalized communities, especially in the Global South (Walsham, 2017; Dell and Kumar, 2016; Avgerou, 2010). Second, while past work often focuses on investigating perspectives of project teams—typically AI developers and nonprofit organizations, emerging research highlights funding agendas’ role in shaping interdisciplinary partnerships (Lin et al., 2024; Saha et al., 2022a; Erete et al., 2023). Through interviews with staff of community organizations participating in AI4SG,  Lin et al. (2024) found that funding agenda influences program goals, proposed solutions, and definitions of success. Focusing on nonprofit organizations, other studies show how funding agendas shape their incentives to adopt emerging tech, and collaboration choices (Bopp et al., 2017; Erete et al., 2016; Goldenfein and Mann, 2023). AI developers and researchers’ approaches to AI ethics and considerations of unintended consequences are also shown to be impacted by funding timeline (Do et al., 2023; Widder et al., 2023; Moitra et al., 2022).
Emphasizing the need to examine how power structures impact technology development in social contexts, scholars have taken up Laura Nader’s (1972) call to “study up” by focusing on systems and actors that shape social systems and exert normative influence (Nader, 1972). Emerging “study up” scholarship in algorithmic fairness (Barabas et al., 2020), public sector AI (Kawakami et al., 2024a), and machine learning data (Miceli et al., 2022) demonstrates that examining those in power can reveal critical leverage points for more accountable AI development. While prior work has highlighted funding’s significant influence on project approaches and outcomes, we still lack a direct understanding of how funding frames AI’s role in addressing social good challenges; there is a need to “study up” AI4SG funding.